OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. 2x + y = −4 y = 3x + 2 (-6/5, -8/5 ) (-8/5, -6/5) (-5/6, -11,5) (-11/5, -6/5)

OpenStudy (texaschic101):

easiest way is the substitution method so we are given : y = 3x + 2 so we are gonna sub in 3x + 2 for y in the other equation 2x + y = -4 2x + 3x + 2 = -4 --- subtract 2 from both sides 2x + 3x = -4 - 2 -- combine like terms 5x = -6 x = -6/5 now, to find y, sub in -6/5 for x in either of the equations y = 3x + 2 y = 3(-6/5) + 2 y = -18/5 + 2 --- common denominator is 5 y = -18/5 + 10/5 y = -8/5 so your solution is : (-6/5, -8/5) any questions at all ??

OpenStudy (anonymous):

No question. Thank you I think I understand it now. Can you help me with one more please ?

OpenStudy (texaschic101):

yes..of course

OpenStudy (anonymous):

A set of equations is given below: Equation C: y = 6x + 9 Equation D: y = 6x + 2 Which of the following options is true about the solution to the given set of equations? One solution No solution Two solutions Infinite solutions

OpenStudy (texaschic101):

set them equal to each other 6x + 9 = 6x + 2 --- subtract 6x from both sides 6x - 6x + 9 = 6x - 6x + 2 -- combine like terms 9 = 2 (not true) NO SOLUTIONS

OpenStudy (anonymous):

thank you so much

OpenStudy (texaschic101):

anytime :)